A distance of 0.002 meter separates two objects of equal mass. If the gravitational force between them is 0.0104N, find the mass of each object.

1
Expert's answer
2022-01-05T14:30:04-0500
"F=\dfrac{Gm^2}{d^2},""m_1=m_2=\sqrt{\dfrac{Fd^2}{G}},""m_1=m_2=\sqrt{\dfrac{0.0104\ N\times(0.002\ m)^2}{6.67\times10^{-11}\ \dfrac{N\times m^2}{kg^2}}}=25\ kg."